OpenStudy (anonymous):

what is -1^2 cuz some people say 1 and others say -1

jimthompson5910 (jim_thompson5910):

\[\Large -1^2 = -1\] but \[\Large (-1)^2 = 1\]

jimthompson5910 (jim_thompson5910):

So that's why most people get confused

OpenStudy (anonymous):

that makes no sense at all

jimthompson5910 (jim_thompson5910):

\[\Large -1^2 = -1*1 = -1\]

jimthompson5910 (jim_thompson5910):

The use of parenthesis changes how we group and multiply, which changes the answer

OpenStudy (anonymous):

 do uk how to solve this